OpenStudy (anonymous):

What is the square root of 90 times the square root of 40?

OpenStudy (anonymous):

√(90)(√(40))=60

OpenStudy (anonymous):

How would you factor that problem out?

OpenStudy (anonymous):

I have answered your previous post with solution

OpenStudy (anonymous):

\[\sqrt{90}*\sqrt{40} = \sqrt{90*40} = \sqrt{3600} = \sqrt{36 * 100} \]\[= \sqrt{36} * \sqrt{100} = 6 * 10\]

OpenStudy (anonymous):

\[\sqrt{90}\times \sqrt{40}\]

OpenStudy (anonymous):

polpak plz ....if u can help me.......c ma quesn

OpenStudy (anonymous):

How would you do this problem: square root of 3x times the square root of 8x to third power?

OpenStudy (anonymous):

The same way. Multiply the stuff under the radical, then pull out squares.

OpenStudy (anonymous):

Start by multiplying the things under the radical, what do you get? \[\sqrt{3x}*\sqrt{8x^3} = \sqrt{???}\]

OpenStudy (anonymous):

24xto the 4th

OpenStudy (anonymous):

Right. So \(x^4 = (?)^2\)